OpenStudy (anonymous):

2.5n^2+5n-17.1 factor that?

OpenStudy (anonymous):

first factor out a 0.1 from the polynomial to get\[=0.1(25x^2+50x-171)\]next find factors of 25*(-171) that sum to 50 and rewrite\[=0.1(25x^2-45x+95x-171)\]factor by grouping\[=0.1\left[ (5x(5x-9)+19(5x-9) \right]\]\[=0.1(5x-9)(5x+19)\]

OpenStudy (anonymous):

awesome thanks